# The Less Intelligent the Elements, the More Intelligent the Whole. Or, Possibly Not?

### v4

@@ Abstract
We approach the debate on how “intelligent” artificial agents should be, by endowing the preys and predators of the Lotka-Volterra model with behavioural algorithms characterized by different levels of sophistication. We find that by endowing both preys and predators with the capability of making predictions based on linear extrapolation a novel sort of dynamic equilibrium appears, where both species co-exist while both populations grow indefinitely. While we confirm that, in general, simple agents favour the emergence of complex collective behaviour, we also suggest that the capability of individuals to take first-order derivatives of one other’s behaviour may allow the collective computation of derivatives of any order.
Keywords: Swarm Intelligence, Collective Intelligence, Crowd Wisdom, Social Connectionism, Prey-Predator, Unlimited Growth
@@ 1 Introduction
Agent-Based Models (ABMs) have been introduced into the social sciences with the aim of reproducing the emergence of macroscopic regularities out of microscopic interactions between relatively simple artificial agents [26], a promise that has obtained some success amid slow but steadily growing diffusion of ABMs in economics, sociology, political science and etology [54]. This paradigm has conceived ABMs as sufficiency proofs that would identify the minimalist microscopic conditions for specific macroscopic dynamics to emerge [25], an attitude succintly described by the KISS principle (Keep It Simple, Stupid) [22]. This posture resonates with the connectionist philosophy of obtaining complex behaviour out of interactions between relatively simple elements, to be applied this time to interacting humans instead of interacting neurons [93] [18].
@@
We approached this question from a slightly different point of view. We focused on one single model, namely an ABM version of the Lotka-Volterra prey-predator model [52] [90]. The reason to select this model is that, beyond obvious ecological interpretations, it is applied to social dynamics as well [16] [94] [101] [17] [59] [40] [92]. Within its framework we explored forms of individual intelligence that have been suggested by empirical findings, theoretical statements and numerical simulations in a wide array of disciplines to be possibly conducive to collective intelligence.
The logic of our method mirrors the practice of exploring the collective consequences of individual behavior in the simulated and iterated multi-player Prisoner’s Dilemma [6] [56]. Just like researchers ask what individual behaviors generate cooperation in the Prisoner’s Dilemma, we ask what individual behaviors generate co-existence of predators and preys in the Lotka-Volterra model.
The rest of this article is organized as follows. In the ensuing section § 2 we distill propositions from several disciplines concerning the relations between individual and collective intelligence. In § 3 we translate these propositions into behavioral algorithms that we ascribe to either predators or preys. Unexpectedly, besides behaviors that generate co-existence of stable populations we also found behaviors that generate co-existence of exploding populations of both predators and preys, a circumstance that we discuss in the concluding section § 4 as expressing unlimited growth in capitalistic economies. The mathematical properties of the Lotka-Volterra model, the parameters and outputs of our model and a sensitivity analysis of its results are expounded in appendices § A, § B and § C, respectively. The code of our extended Lotka-Volterra model is available on CoMSES.
@@ 2 Individual and Collective Intelligence
In this section we review theoretical insights and empirical evidence that bear on the question whether a collectivity works better if their members understand its underlying mechanisms, or rather base their decisions on simple rules. We subsumed our findings into three propositions, the third of which comes in three versions. We list these propositions in ascending order, from KISS- to increasingly KIDS-supportive.
Furthermore, we also collected three technical qualifications that may be relevant to these propositions. These technical qualifications are expounded in § 2.1.
Our first proposition captures positions that justify the KISS principle on theoretical and empirical reasons. According to their proponents, KISS is not just a convenient and sometimes sufficient assumption but also what humans actually do (descriptive value), as well as what they should do (normative value).
One extreme version maintains that individual intelligence is totally irrelevant for social science [65]. Inparticular, several “zero-intelligence” models can be found in economics, notably for double-auction markets [33] [34], financial markets [23] and, more in general, any sort of prediction-based markets [67]. Such models typically suggest that markets — even futures markets — can work perfectly well with decision makers who have no memory and are incapable of cognition of whatsoever form. However, some of these models have been found to hide constraints that artificially generate their most impressive results [32].
@@
The FFH paradigm does not understand bounded rationality [76] [77] as a deviation from optimality but rather as the evolutionary optimal answer to radical uncertainty and imperfect information [71] [39] [30] [63]. Consequently, FFH have descriptive as well as prescriptive value [29] [43] [37] [19].
Similar conclusions are suggested by The Traveler’s Dilemma, a game where rational players endowed with unlimited ability to read one other’s minds (“I think that she thinks that I think that …”) receive lower payoffs than boundedly rational players who carry out mind reading for a couple of steps at most [8] [9]. Interestingly, when this game is played with real people they invariably behave the “stupid,” more efficient way.
Finally, ABMs of the stock market where both sophisticated and simple agents operate have generated the unexpected finding that sophisticated agents typically make lower profits than the simple ones [82] [36]. In this specific case, the reason was that sophisticated estimation techniques suffered from overfitting a noisy environment. By contrast, simple rules traded-off their imperfections with their robustness.
Let us condense these KISS-supportive experiences into the following proposition:
@@
Let us now move onto a set of insights that focus on a very special aspect of individual intelligence, namely the ability to make predictions. Prediction can be very local and very limited, such as linear extrapolation, or it may involve extremely sophisticated cognitive abilities. Depending on flavour, it can span the whole gamut from moderate KISS to maximum KIDS. In this section we limit ourselves to enunciate a general principle that we shall decline into specific choices in § 3.
Several ABMs have highlighted the importance of predictions. For instance, it has been found that robots that are capable of mapping their environment and predicting what other robots will eventually do perform better than those that are based on FFH [73]. Likewise, models of distributed resource allocation have highlighted that agents that are able to extrapolate general trends perform better than those who do not [45] [46].
Shifting from simulations to the real world, human groups act as teams insofar their members make the effort of reading one other’s mind, predict what their mates will do, and behave accordingly [47] [91]. Experimental tests on indicators of social attentiveness, such as eye movements or taking turns at speaking, confirm this point [100]. Predictions in teams clearly go beyond mere extrapolation, but still, they are not as complex as having an overall picture of how the whole team is working. The coach of a sporting team may have it, but all what is required by team players is that they observe one another, predict, and coordinate.
Out of these considerations we distill the notion that individual ability to make predictions can contribute to collective intelligence. We subsume this concept by means of the following proposition:
@@
Finally, let us consider maximum KIDS. While one can think of several fields other than prediction, the ability to think and understand collective behaviour stands out as a clear benchmark for the maximum sophistication of individual agents. This idea can be traced back to Ross Ashby, who called attention onto the difference between the dynamics expressed by the variables of a system and those induced by changing its parameters [5] [84]. At least three influential authors such as Gregory Bateson, Chris Argyris and Donald Schön have been inspired by Ashby when pointing to the importance of understanding system-wide dynamics ([10], p. 292, [3], p. 21, footnote n. 1) [87].
Henceforth, we shall express this idea in terms of Lyapunov functions, saying that while short-sighted decision-makers can be represented by means of variables descending along given basins of attraction, those who are capable to discern collective dynamics are also contemplating the possibility of changing the shape of those basins (see § 3).
Following customary usage in Psychology, let us employ the term meta-cognition for an individual’s capability to understand collective dynamics:
@@
One notable feature of Proposition 3 is that it does not immediately translate into specific consequences for the collective. One critical reason is that, given an individual who is capable of meta-cognition, his possibilities for affecting the organization depend on hierarchical position [66] [49]. Specifically, we singled out three versions of Proposition 3 that differ from one another in this respect.
@@
By contrast, if those individuals who envisage the collective intelligence of their organization do not have the authority to steer it, deutero-learning ensues [10] [11]. This concept stems from anthropology and it rather fits the case of individuals who must passively accept a mechanism of which they become aware, whose aims they do not share [86] [88]. Contradiction between official declarations and real experience generates conflicting obligations known as double bind [11] [74]. This point of view is entailed in the following version of Proposition 3:
@@
Finally, a still different case occurs when the collective intelligence of an organization is envisaged by the vast majority of its stakeholders who openly agree on a course of action [48] [99]. The third version of Proposition 3 ensues:
@@ 2.1 Technical Qualifications
The emergence of collective dynamics out of individual behaviour has also technical features that are unrelated to the degree of individual intelligence. These technical features can eventually qualify the above propositions.
The first qualification requires individuals to be heterogeneous in order to generate interesting collective dynamics. This qualification applies even to extreme KISS approaches such as artificial neural networks, whose neurons are as simple as summators and yet require random initialization of weights. In more complex models heterogeneity increases the range where self-organized critical states perform the most complex computations [72]. One intuitive confirmation comes from the electrical signals of real brains, whose neurons fire in unison under epileptic seizures but exhibit complex patterns otherwise [12] [38].
Group psychoanalysis offers a similar insight concerning mass behaviour, which can slip into irrational modes where individuals merge their minds generating collectives that do not exhibit any intelligence at all (e.g., certain religious sects, or crowds following populist leaders) [14] [81]. By contrast, crowd wisdom — the ability of large groups to provide solutions that would be difficult for isolated individuals to reach [44] — relies on heterogeneity of individual interpretations [41] [42].
The above considerations can be subsumed in the following qualification:
@@
The second technical qualification concerns collective agreements and it is obviously most relevant for Proposition § 3.3. Collective agreements generate stable equilibria, but individual intelligence can still make a difference isofar it concerns the speed of convergence.
Macroeconomics has a good point in case. In the 1970s, many Central Banks were trying to stimulate the economy by increasing the money supply in the belief that economic actors would use it to increase GDP, but those actors were eventually able to anticipate that inflation would be the long-term consequence, which expectation generated inflation already in the short run [53]. In this case, greater individual intelligence — labeled as rational expectations — made for faster convergence to a high-inflation, high-unemployment equilibrium. Note that in this case the “agreement” exists only among actors other than the Central Bank (which is treated as an exogenous disturbance, indeed).
Let us capture this insight by means of the following qualification:
@@
Finally, the third technical qualification concerns the possibility that small individual decisions have a sizeable impact on collectives. Specifically, the butterfly effect states that non-linearities, through chains of bifurcations, are capable of generating chaotic dynamics which may revolve around a “strange attractor” nevertheless [51]. We need a sort of generalized-butterfly qualification that is not limited to chaotic dynamics but includes any sizeable change of the Lyapunov function as well. Let us express this extended butterfly effect by means of the following qualification:
@@ 3 Predators and Preys
In this section we apply the aforementioned insights to the Lotka-Volterra model of prey-predator dynamics [52] [90]. Specifically, we explore whether any sort of collective intelligence appears by endowing predators and preys with variants of individual intelligence inspired by the propositions and qualifications outlined in § 2.
In general, co-existence of predators and preys is the preferred outcome in any application of the Lotka-Volterra model, either because one desires some sort of ecological equilibrium, or because the model performs more sophisticated, more interesting computations if species co-exist, or both. Henceforth, we shall take species co-existence as an indication of collective intelligence, looking for sorts of individual intelligence that make it appear.
There are several reasons for focusing on the Lotka-Volterra model. The first one is that it is relevant to a number of settings other than animal species, including as diverse applications as the business cycle [35] [55] [78] [101], technological substitution [13] [62] [101], ideological and power struggles [89] [17] [59] [40] and market-share dynamics [94] [58] [92].
Thus, it arguably reflects important aspects of human relations and interactions. Exploring the possibility that predators and preys are capable of sophisticated forms intelligence reflects these sorts of applications rather than the purely biological ones, whereas simpler algorithms are likely to be expressed by less sophisticated brains.
@@
However, existing ABMs achieve consistency with the aggregate Lotka-Volterra by adding features that are specific to animal populations, such as availability of food and shelter for preys, or stages of phenotypic development for predators [97] [98] [96] [15] [83]. By contrast, we are interested in collective dynamics that can be influenced by sophisticated individual decisions. Thus, a final reason is that this endeavour has not been attempted hitherto.
Let us begin with the basic Lotka-Volterra model. Let $x\in\Re^{+}_{0}$ and $y\in\Re^{+}_{0}$ denote the environmental density  [^1] By “environmental density” we mean the ratio of the number of individuals to some measure of the size of the natural environment where predators and preys live, such as its area or volume. of preys and predators, respectively. In its simplest version, the Lotka-Volterra model captures their dynamics by means of the following pair of differential equations:
@@
Let us turn to the discrete-time, agent-based version of the Lotka-Volterra model in order to analyze what behavioral algorithms take the system into either $E_{1}$, $E_{2}$ or $E_{3}$, respectively. Notably, moving onto computational models based on heterogeneous interacting agents implies that Technical Qualification 1 applies.
Henceforth we shall start from Wilensky and Reisman’s Wolf Sheep Predation model [95] [96] on the NetLogo platform where this model is available in the standard library. In order to experiment with behavioral hypotheses we added additional features to the basic model. Our extended version is available on CoMSES.
In Wilensky and Reisman’s Lotka-Volterra, $E_{3}$ almost never sustains itself. Specifically, we observed $E_{1}$ (both preys and predators disappear) $41\%$ of the times, $E_{2}$ (only preys survive) $58.7\%$ of the times whereas only $0.3\%$ of the times the two species managed to co-exist (see Appendix B for details). Wilensky and Reisman obtained stable cycles by adding a renewable source of energy for preys [96].
We took on a different route. We started from the basic, unstable model that almost invariably ends into either $E_{1}$ or $E_{2}$, and experimented with predators’ and preys’ intelligence in order to discover whether either predators’ or preys’ intelligence, or both, can modify its Lyapunov function to make $E_{3}$ the center of a new basin of attraction.
Figure 1 illustrates this concept by means of fictional Lyapunov functions. On the left (a), a saddle from which the system sooner or later ends either in $E_{1}$ or in $E_{2}$. This is the sort of Lyapunov function that likely describes the Wilensky-Reisman model as it is. We would like that preys and predators behave in ways such that the Lyapunov function becomes either as in (b), in which case both species coexist with oscillating populations, or (c), in which case constant populations of predators and preys obtain. Our research question is whether any sort of individual intelligence exists, that helps them reach one of these configurations.
@@
Henceforth we explore what happens to Wilensky and Reisman’s model by endowing preys and predators with alternative sorts of intelligence inspired by the Propositions and Qualifications expounded in § 2 and § 2.1, respectively. Details about the model are in Appendix B. All reported results have been averaged over 1,000 runs.

### v5

@@ Abstract
The agent-based modelling community has a debate on how “intelligent” artificial agents should be, and in what ways their local intelligence relates to the emergence of a collective intelligence. I approach this debate by endowing the preys and predators of the Lotka-Volterra model with behavioral algorithms characterized by different levels of sophistication. The main finding is that by endowing both preys and predators with the capability of making predictions based on linear extrapolation a novel sort of dynamic equilibrium appears, where both species co-exist while both populations grow indefinitely. While this broadly confirms that, in general, relatively simple agents favor the emergence of complex collective behavior, it also suggests that one fundamental mechanism is that the capability of individuals to take first-order derivatives of one other’s behavior can allow the collective computation of derivatives of any order.
Keywords: Collective Intelligence, Swarm Cognition, Crowd Wisdom, Social Connectionism, Prey-Predator, Unlimited Growth
@@ 1 Introduction
Agent-Based Models (ABMs) have been introduced into the social sciences with the aim of reproducing the emergence of macroscopic regularities out of microscopic interactions between relatively simple artificial agents [25], a promise that has obtained some success amid slow but steadily growing diffusion of ABMs in economics, sociology, political science and ethology [53]. This paradigm has conceived ABMs as sufficiency proofs that would identify the minimalist microscopic conditions for specific macroscopic dynamics to emerge [24], an attitude succinctly described by the KISS principle (Keep It Simple, Stupid) [21]. This posture resonates with the connectionist philosophy of obtaining complex behavior out of interactions between relatively simple elements, to be applied this time to interacting (simulated) humans instead of interacting (simulated) neurons [93] [17] [84].
@@
We approached this question from a slightly different point of view. We focused on one single model, namely an ABM version of the Lotka-Volterra prey-predator model [51] [90]. The reason to select this model is that, beyond obvious ecological interpretations, the Lotka-Volterra is applied to social dynamics as well [15] [94] [101] [16] [58] [39] [92]. Within its framework I explored forms of individual intelligence that have been suggested by empirical findings, theoretical statements and numerical simulations in a wide array of disciplines to be possibly conducive to collective intelligence.
The logic of this method mirrors the practice of exploring the collective consequences of individual behavior in the simulated and iterated multi-player Prisoner’s Dilemma [5] [55]. Just like other researchers ask what individual behaviors generate cooperation in the Prisoner’s Dilemma, I ask what individual behaviors generate co-existence of predators and preys in the Lotka-Volterra model.
The rest of this article is organized as follows. In the ensuing section § 2 I distill propositions from several disciplines concerning the relations between individual and collective intelligence. In § 3 I translate these propositions into behavioral algorithms that can be ascribed to either predators, or preys. Unexpectedly, besides behaviors that generate co-existence of stable populations there exist behavior algorithms that generate co-existence of exploding populations of both predators and preys, a circumstance that I discuss in the concluding section § 4 as remindful of unlimited growth in capitalistic economies. The mathematical properties of the Lotka-Volterra model, the parameters and outputs of our model and a sensitivity analysis of its results are expounded in appendices § A, § B and § C, respectively.  [^1] The code of the extended Lotka-Volterra model is available on CoMSES. This code has been written by Andrea Policarpi upon inspiration from Wilensky and Reisman’s Wolf Sheep Predation model [95] [96], available among the library models of the NetLogo platform.
@@ 2 Individual and Collective Intelligence
In this section I review theoretical insights and empirical evidence that bear on the question whether a collectivity works better if their members understand its underlying mechanisms, or rather enact simple rules. This survey wants to be as broadly interdisciplinary as possible, ranging from the computational to the social sciences and beyond, to the neurosciences, group psychology and psychoanalisis. Its findings are essential in order to carry out a reasoned, rather than random exploration of behavioral algorithms for predators and preys.
Henceforth, the results of this survey are condensed in three propositions, the third of which comes in three versions. These propositions are listed in ascending order, from KISS- to increasingly KIDS-supportive. Propositions are complemented by three technical qualifications expounded in § 2.1.
The first proposition subsumes positions that justify the KISS principle on theoretical and empirical grounds. Its implicit assumption is that KISS is not just a convenient and sometimes sufficient assumption but also a reliable guideline to describe what humans actually do (descriptive value), as well as what they should do (normative value).
One extreme version maintains that individual intelligence is totally irrelevant for social science [64]. Specifically, several economic “zero-intelligence” models exist for double-auction markets [32] [33], financial markets [22] and, more in general, any sort of prediction-based markets [66]. Such models typically suggest that markets — even futures markets — can work perfectly well with decision makers who have no memory and are incapable of cognition whatsoever. However, some of these models have been found to hide constraints that artificially generate their most impressive results [31].
@@
Notably, the FFH paradigm does not understand bounded rationality [76] [77] as a deviation from optimality but rather as the evolutionary optimal answer to radical uncertainty and imperfect information [70] [38] [29] [62]. Hence, its descriptive and prescriptive value [28] [42] [36] [18].
Similar conclusions are suggested by The Traveler’s Dilemma, a game where rational players endowed with unlimited ability to read one other’s minds (i.e., “I think that she thinks that I think that …”) receive lower payoffs than boundedly rational players who carry out mind reading for a couple of steps at most [7] [8]. Interestingly, when this game is played with real people they invariably behave the “stupid,” more efficient way.
Finally, ABMs of the stock market where agents with diverse degrees of “intelligence” operate have generated the unexpected finding that sophisticated agents typically make lower profits than the simpler ones [82] [35]. In this specific case, the reason was that sophisticated estimation techniques suffered from over-fitting a noisy environment; by contrast, simple rules traded-off imperfections with robustness.
The following proposition is able to condense the above KISS-supportive stances in a sufficiently strong (prescriptive) though not absolute manner:
@@
In the aforementioned survey, the ability to make predictions surfaced quite often as an aspect of individual intelligence that has a substantial impact on collective intelligence. Prediction can be very local and very limited, such as linear extrapolation, or it may involve extremely sophisticated cognitive abilities. Correspondingly, it can span the whole gamut from moderate KISS to maximum KIDS.
Several simulations have highlighted the importance of predictions. For instance, it has been found that robots that are capable of predicting the behavior of other robots can achieve better performances [73]; likewise, in ABMs of distributed resource allocation agents that are able to extrapolate general trends perform better than those who do not [44] [45].
Shifting from simulations to the real world, human groups act as teams insofar their members make the effort of reading one other’s mind, predict what their mates will do, and behave accordingly [46] [91]. Anecdotical evidence is confirmed by experimental tests on indicators of social attentiveness, such as eye movements or taking turns at speaking [100].
Out of these considerations one can distill the notion that individual ability to make predictions can contribute to collective intelligence. By expressing a potentiality, the following proposition can accomodate several stances in the KISS–KIDS gamut:
@@
Finally, maximum KIDS can be approached by taking the ability to think and understand collective behavior as the maximum sophistication of individual agents [9, p. 292] [3, p. 21] [87]. In psychology and elsewhere, this ability is known as meta-cognition. In systems-theoretical terms, the behavior of short-sighted decision-makers can be subsumed by variables descending along given basins of attraction of a Lyapunov function, whereas agents who are capable to discern collective dynamics can also contemplate the possibility of changing the shape of those basins (see § 3).
Since maximum KIDS is highly hypothetical, the corresponding proposition must be hypothetical as well. The common core of several statements to be found in the literature can be expressed as follows:
@@
One notable feature of Proposition 3 is that it does not immediately translate into specific consequences for the collective. One reason is that, given an individual who is capable of meta-cognition, their possibilities for affecting the organization depend on hierarchical position [65] [48]. Specifically, Proposition 3 can be found in three versions that differ from one another in this respect.
@@
By contrast, if those individuals who envisage the collective intelligence of their organization do not have the authority to steer it, deutero-learning ensues [9] [10]. This concept stems from anthropology and it rather fits the case of individuals who must passively accept a mechanism of which they become aware, but whose aims they do not share [86] [88]. Contradiction between official declarations and real experience generates conflicting obligations known as double bind [10] [74]. This point of view can be expressed by the following version of Proposition 3:
@@
Finally, a still different case occurs when the collective intelligence of an organization is envisioned by the vast majority of its stakeholders, who openly agree on a course of action [47] [99]. This third version of Proposition 3 can be summarised as follows:
@@ 2.1 Technical Qualifications
The emergence of collective dynamics out of individual behavior has also technical features that are unrelated to the degree of individual intelligence. These technical features can eventually qualify the above propositions.
The first qualification requires individuals to be heterogeneous in order to generate interesting collective dynamics. This qualification applies even to extreme KISS approaches such as artificial neural networks, whose neurons are as simple as summators and yet require random initialization of their weights. In more complex models, heterogeneity increases the range of computations [71] [72] (one example is provided by the electrical signals generated by biological neurons, which fire in unison under epileptic seizures but exhibit complex patterns otherwise [11] [37]; in this case, heterogeneity is essential for intelligence).
Moving on to collective behavior, group psychoanalysis identifies irrational collective behavior modes characterized by individuals merging their minds into a homogeneous, not-so-intelligent whole as, e.g., in religious or political sects [13] [81]. By contrast, crowd wisdom — the ability of large groups to provide solutions that would be difficult for isolated individuals to reach [43] — relies on heterogeneity of individual interpretations [40] [41].
The above considerations can be subsumed in the following, rather simple qualification:
@@
The second qualification concerns collective agreements. Thus, it is most relevant for Proposition § 3.3.
This qualification makes the point that although collective agreements necessarily generate stable equilibria, individual intelligence can still make a difference isofar it concerns the speed of convergence. In some contexts, speed makes the difference.
Macroeconomics has a good point in case. In the 1970s, many Central Banks were trying to stimulate the economy by increasing the money supply in the belief that economic actors would use it to increase GDP. However, those actors were eventually able to anticipate that inflation would be the long-term consequence, which expectation would quickly generate inflation [52]. In this case, greater individual intelligence — known in economics as rational expectations — made for faster convergence to a high-inflation, high-unemployment equilibrium (note that in this case the “agreement” exists only among actors other than the Central Bank, which is treated as an exogenous disturbance, indeed).
This insight by means of the following qualification:
@@
Finally, the third technical qualification concerns the possibility that small individual decisions have a sizeable impact on collectives. Specifically, the butterfly effect states that non-linearities, through chains of bifurcations, are capable of generating chaotic dynamics which may revolve around a “strange attractor” nevertheless [50]. This qualification can be expressed as follows:
@@ 3 Predators and Preys
This section endows the predators and preys of the Lotka-Volterra model [51] [90] with variants of individual intelligence inspired by the propositions and qualifications outlined in § 2. In general, co-existence of predators and preys is the preferred outcome, either because one desires some sort of ecological equilibrium, or because the model performs more sophisticated, more interesting computations if species co-exist, or both. Thus, species co-existence can be taken as an indication of collective intelligence. Consequently, this section pursues the goal of identifying individual behavioral algorithms that make preys and predators co-exist.
Species co-existence is an attractive instance of collective intelligence for at least two reasons. The first one is that the Lotka-Volterra model is relevant to a number of settings other than animal species, including as diverse applications as the business cycle [34] [54] [78] [101], technological substitution [12] [61] [101], ideological and power struggles [89] [16] [58] [39] and market-share dynamics [94] [57] [92].
Thus, it arguably reflects important aspects of human relations and interactions. Exploring the possibility that predators and preys are capable of sophisticated forms of intelligence reflects these sorts of applications, rather than the purely biological ones.
@@
However, existing ABMs achieve consistency with the aggregate Lotka-Volterra by adding features that are specific to animal populations, such as availability of food and shelter for preys, or stages of phenotypic development for predators [97] [98] [96] [14] [83]; by contrast, we are interested in achieving species co-existence out of (sophisticated individual decisions).
Let us first of all introduce the basic Lotka-Volterra model. Let $x\in\Re^{+}_{0}$ and $y\in\Re^{+}_{0}$ denote the environmental densities  [^2] Environmental density is the ratio of the number of individuals to some measure of the size of the natural environment where predators and preys live, such as its area or volume. of preys and predators, respectively. In its simplest version, the Lotka-Volterra model captures their dynamics by means of the following pair of differential equations:
@@
Let us turn to the discrete-time, agent-based version of the Lotka-Volterra model in order to analyze what behavioral algorithms take the system into either $E_{1}$, $E_{2}$ or $E_{3}$, respectively. Notably, moving onto computational models based on heterogeneous interacting agents implies that Technical Qualification 1 applies. In particular, I move from Wilensky and Reisman’s Wolf Sheep Predation model [95] [96]  [^3] Available in NetLogo, models’ library. to modify its behavioral algorithms in subsequent steps  [^4] This extended version is available at CoMSES.
In Wilensky and Reisman’s Lotka-Volterra, $E_{3}$ almost never sustains itself. Specifically, $E_{1}$ (both preys and predators disappear) was observed $41\%$ of the times, $E_{2}$ (only preys survive) $58.7\%$ of the times whereas only $0.3\%$ of the times the two species managed to co-exist (see Appendix B for details). Indeed, Wilensky and Reisman had to add a renewable source of energy for preys in order to obtain stable cycles [96].
I took on a different route. I started from the basic, unstable model that almost invariably ends into either $E_{1}$ or $E_{2}$, and experimented with predators’ and preys’ intelligence in order to discover whether any of them can modify its Lyapunov function to make $E_{3}$ the center of a new basin of attraction.
Figure 1 illustrates this concept by means of fictional Lyapunov functions. On the left (a), a saddle from which the system sooner or later ends either in $E_{1}$ or in $E_{2}$. This is the sort of Lyapunov function that likely describes the Wilensky-Reisman model as it is. We rather want that the Lyapunov function is either as in (b), in which case both species coexist with oscillating populations, or (c), in which case constant populations of predators and preys obtain. The research question is whether any sort of individual intelligence exists, that helps them reach one of these configurations.
@@
Henceforth, the results of modified versions of Wilensky and Reisman’s model where preys and predators are endowed with behavioral algorithms inspired by the Propositions and Qualifications expounded in § 2. Details about the model are in Appendix B. All reported results have been averaged over 1,000 runs.